# Probing Green’s Function Zeros by Co-tunneling through Mott Insulators

### v1

@@ A.3 First order perturbation theory
$M$ $=\sum_{k,\sigma}\mu_{k,\sigma}n_{k,\sigma}=\sum_{k}\frac{1}{2}(\mu_{k,\uparrow}+\mu_{k,\downarrow})(n_{k,\uparrow}+n_{k,\downarrow})+\frac{1}{2}(\mu_{k,\uparrow}-\mu_{k,\downarrow})(n_{k,\uparrow}-n_{k,\downarrow})$
$=\sum_{k}\bar{\mu}_{k}(n_{k,\uparrow}+n_{k,\downarrow})+h_{k}\frac{1}{2}(n_{k,\uparrow}-n_{k,\downarrow})=\sum_{k}\bar{\mu}_{k}(n_{k,\uparrow}+n_{k,\downarrow})+h_{k}S^{z}_{k}\text{.}$ (43)
Here, we decomposed the terms into an spin-isotropic potential $\bar{\mu}_{k}$ and a local magnetic field $h_{k}$. The individual components of $G^{1}$ (compare to Eqs.(40), (41),(42)) can be simplified as:
$\Delta_{0}$ $=\langle M\rangle=\sum_{k}\bar{\mu}_{k}+\sum_{k}h_{k}\langle S^{z}_{k}\rangle$ (44)
@@
$R_{j,\sigma;j^{\prime},\sigma^{\prime}}$ $=\langle c_{j,\sigma}(M-\Delta_{0})c_{j^{\prime},\sigma^{\prime}}^{\dagger}\rangle$
$=\sum_{k}\bar{\mu}_{k}\langle c_{j,\sigma}\left[\left(n_{k,\uparrow}+n_{k,\downarrow}\right)-1\right]c_{j^{\prime},\sigma^{\prime}}^{\dagger}\rangle+\sum_{k}h_{k}\langle c_{j,\sigma}\left[S^{z}_{k}-\langle S^{z}_{k}\rangle\right]c_{j^{\prime},\sigma^{\prime}}^{\dagger}\rangle$
$=\sum_{k}\bar{\mu}_{k}\langle c_{j,\sigma}\left[n_{k,\uparrow}+n_{k,\downarrow},c_{j^{\prime},\sigma^{\prime}}^{\dagger}\right]\rangle$
$+\frac{1}{2}\sum_{k}h_{k}\langle c_{j,\sigma}c_{j^{\prime},\sigma^{\prime}}^{\dagger}\left[S^{z}_{k}-\langle S^{z}_{k}\rangle\right]\rangle+\frac{1}{2}\sum_{k}h_{k}\langle\left[S^{z}_{k}-\langle S^{z}_{k}\rangle\right]c_{j,\sigma}c_{j^{\prime},\sigma^{\prime}}^{\dagger}\rangle$
$+\frac{1}{2}\sum_{k}h_{k}\langle c_{j,\sigma}\left[S^{z}_{k},c_{j^{\prime},\sigma^{\prime}}^{\dagger}\right]\rangle+\frac{1}{2}\sum_{k}h_{k}\langle\left[c_{j,\sigma},S^{z}_{k}\right]c_{j^{\prime},\sigma^{\prime}}^{\dagger}\rangle$
$=\bar{\mu}_{j}\delta_{j,j^{\prime}}\left(\frac{1}{2}\tau_{0}-\overrightarrow{\tau}\cdot\langle\overrightarrow{S}_{j}\rangle\right)_{\sigma,\sigma^{\prime}}+\frac{1}{2}\delta_{j,j^{\prime}}\langle\left[\left(\frac{1}{2}\tau_{0}-\overrightarrow{\tau}\cdot\overrightarrow{S}_{j}\right)_{\sigma,\sigma^{\prime}},\sum_{k}h_{k}\left[S^{z}_{k}-\langle S^{z}_{k}\rangle\right]\right]_{+}\rangle$
$+\frac{h_{j}}{4}\delta_{j,j^{\prime}}\left[\left(\frac{1}{2}\tau_{0}-\overrightarrow{\tau}\cdot\langle\overrightarrow{S}_{j}\rangle\right)_{\sigma,\sigma^{\prime}},\sigma_{z}\right]_{+}$
$=\delta_{j,j^{\prime}}\frac{1}{2}\left(\bar{\mu}_{j}\tau_{0}+\frac{h_{j}}{2}\sigma_{z}\right)_{\sigma,\sigma^{\prime}}-\delta_{j,j^{\prime}}\left(\frac{h_{j}}{2}\langle S^{z}_{j}\rangle\tau_{0}+\bar{\mu}_{j}\overrightarrow{\tau}\cdot\langle\overrightarrow{S}_{j}\rangle\right)_{\sigma,\sigma^{\prime}}$
$+\frac{1}{2}\delta_{j,j^{\prime}}\langle\left[\left(\frac{1}{2}\tau_{0}-\overrightarrow{\tau}\cdot\overrightarrow{S}_{j}\right),\sum_{k}h_{k}(S^{z}_{k}-\langle S^{z}_{k}\rangle)_{\sigma,\sigma^{\prime}}\right]_{+}\rangle$ (46)
$V_{j,\sigma;j^{\prime},\sigma^{\prime}}$ $=\langle c_{j^{\prime},\sigma^{\prime}}^{\dagger}(M-\Delta_{0})c_{j,\sigma}\rangle$
$=\sum_{k}\bar{\mu}_{k}\langle c_{j^{\prime},\sigma^{\prime}}^{\dagger}\left[\left(n_{k,\uparrow}+n_{k,\downarrow}\right)-1\right]c_{j,\sigma}\rangle+\sum_{k}h_{k}\langle c_{j^{\prime},\sigma^{\prime}}^{\dagger}\left[S^{z}_{k}-\langle S^{z}_{k}\rangle\right]c_{j,\sigma}\rangle$
$=\sum_{k}\bar{\mu}_{k}\langle c_{j^{\prime},\sigma^{\prime}}^{\dagger}\left[n_{k,\uparrow}+n_{k,\downarrow},c_{j,\sigma}\right]\rangle$
$+\frac{1}{2}\sum_{k}h_{k}\langle c_{j^{\prime},\sigma^{\prime}}^{\dagger}c_{j,\sigma}\left[S^{z}_{k}-\langle S^{z}_{k}\rangle\right]\rangle+\frac{1}{2}\sum_{k}h_{k}\langle\left[S^{z}_{k}-\langle S^{z}_{k}\rangle\right]c_{j^{\prime},\sigma^{\prime}}^{\dagger}c_{j,\sigma}\rangle$
$+\frac{1}{2}\sum_{k}h_{k}\langle c_{j^{\prime},\sigma^{\prime}}^{\dagger}\left[S^{z}_{k},c_{j,\sigma}\right]\rangle+\frac{1}{2}\sum_{k}h_{k}\langle\left[c_{j^{\prime},\sigma^{\prime}}^{\dagger},S^{z}_{k}\right]c_{j,\sigma}\rangle$
$=-\bar{\mu}_{j}\delta_{j,j^{\prime}}\left(\frac{1}{2}\tau_{0}+\overrightarrow{\tau}\cdot\langle\overrightarrow{S}_{j}\rangle\right)_{\sigma,\sigma^{\prime}}+\frac{1}{2}\delta_{j,j^{\prime}}\langle\left[\left(\frac{1}{2}\tau_{0}+\overrightarrow{\tau}\cdot\overrightarrow{S}_{j}\right)_{\sigma,\sigma^{\prime}},\sum_{k}h_{k}\left[S^{z}_{k}-\langle S^{z}_{k}\rangle\right]\right]_{+}\rangle$
$-\frac{h_{j}}{4}\delta_{j,j^{\prime}}\left[\left(\frac{1}{2}\tau_{0}+\overrightarrow{\tau}\cdot\langle\overrightarrow{S}_{j}\rangle\right)_{\sigma,\sigma^{\prime}},\sigma_{z}\right]_{+}$
$=-\delta_{j,j^{\prime}}\frac{1}{2}\left(\bar{\mu}_{j}\tau_{0}+\frac{h_{j}}{2}\sigma_{z}\right)_{\sigma,\sigma^{\prime}}-\delta_{j,j^{\prime}}\left(\frac{h_{j}}{2}\langle S^{z}_{j}\rangle\tau_{0}+\bar{\mu}_{j}\overrightarrow{\tau}\cdot\langle\overrightarrow{S}_{j}\rangle\right)_{\sigma,\sigma^{\prime}}$
$+\frac{1}{2}\delta_{j,j^{\prime}}\langle\left[\left(\frac{1}{2}\tau_{0}+\overrightarrow{\tau}\cdot\overrightarrow{S}_{j}\right)_{\sigma,\sigma^{\prime}},\sum_{k}h_{k}\left[S^{z}_{k}-\langle S^{z}_{k}\rangle\right]\right]_{+}\rangle$ (47)
@@
$G^{1}_{j,\sigma;j^{\prime},\sigma^{\prime}}$ $=\delta_{j,j^{\prime}}\left[\left(\bar{\mu}_{j^{\prime}}\tau_{0}+\frac{h_{j}}{2}\sigma_{z}\right)-\langle\left[\overrightarrow{\tau}\cdot\overrightarrow{S}_{j},\sum_{k}h_{k}(S^{z}_{k}-\langle S^{z}_{k}\rangle)\right]_{+}\rangle\right]_{\sigma,\sigma^{\prime}}$
$=M_{j,\sigma;j^{\prime},\sigma^{\prime}}-\delta_{j,j^{\prime}}\left[\langle\left[\overrightarrow{\tau}\cdot\overrightarrow{S}_{j},\sum_{k}h_{k}(S^{z}_{k}-\langle S^{z}_{k}\rangle)\right]_{+}\rangle\right]_{\sigma,\sigma^{\prime}}$ (48)
Notice that for an $SU(2)$-symmetric bare Hamiltonian $M$ the magnetic field $h_{k}$ vanishes, and $G^{1}$ equals $M$.
As second bare hopping Hamiltonian we investigate a typical spin-isotropic hopping $W$($\equiv H_{0,l=1}$):
@@
The final expression for $G^{1}$ corresponding to the total bare Hamiltonian, $H_{0}=W+M$, is obtained by summing all individual contributions to $G^{1}$. Using $G^{0}$ and $G^{1}$, we can determine the Hamiltonian of zeros, $H^{\text{z}}$. For a spin-isotropic $H_{0}$ (i.e., in the absence of a magnetic field, $h_{j}=0~{}\forall~{}j$) we find:
@@
As a final result, and in line with [33], we observe that the local terms of the bare Hamiltonian $M$ reemerge unrenormalized in the $H^{z}$-Hamiltonian, while the non-local terms $W$ are rescaled by the spin correlation function at the corresponding bond, $\langle\overrightarrow{S}_{j}\cdot\overrightarrow{S}_{j^{\prime}}\rangle_{j\neq j^{\prime}}$. The expectation values are taken with respect to the groundstate $|GS_{0}\rangle$ in the unperturbed limit,i.e., is the groundstate of an effective spin-$1/2$ model (degenerated perturbation theory). Although $|GS_{0}\rangle$ is defined within the subspace of the atomic limit, it remains generally non-trivial, i.e., long-range correlations can occur:
@@
Due to the additive structure of the GF, Eq.(14), our results are valid for any density matrix $\rho=\sum_{\nu}w_{\nu}|E^{P,m}_{\nu,l}\rangle\langle E^{P,m}_{\nu,l}|$, which contains only states from the lowest energy sector $m=0,~{}l=0$. This includes approximately the Gibbs state at very low temperatures. All the expectation values have to be simply updated:
@@
A.5 Self-consistent selfenergy construction

### v2

@@ A.4 Rewriting the GF-approximation using spin operators
$M$ $=\sum_{k}\bar{\mu}_{k}(n_{k,\uparrow}+n_{k,\downarrow})\text{.}$ (43)
Here, we obtain the local spin-isotropic potentials $\bar{\mu}_{k}$. The individual components of $G^{1}$ (compare to Eqs.(40-42)) can be simplified as:
$\Delta_{0}$ $=\langle M\rangle=\sum_{k}\bar{\mu}_{k}$ (44)
@@
$R_{j,\sigma;j^{\prime},\sigma^{\prime}}$ $=\langle c_{j,\sigma}(M-\Delta_{0})c_{j^{\prime},\sigma^{\prime}}^{\dagger}\rangle=\sum_{k}\bar{\mu}_{k}\langle c_{j,\sigma}\left[\left(n_{k,\uparrow}+n_{k,\downarrow}\right)-1\right]c_{j^{\prime},\sigma^{\prime}}^{\dagger}\rangle$
$=\sum_{k}\bar{\mu}_{k}\langle c_{j,\sigma}\left[n_{k,\uparrow}+n_{k,\downarrow},c_{j^{\prime},\sigma^{\prime}}^{\dagger}\right]\rangle=\bar{\mu}_{j}\delta_{j,j^{\prime}}\left(\frac{1}{2}\tau_{0}-\overrightarrow{\tau}\cdot\langle\overrightarrow{S}_{j}\rangle\right)_{\sigma,\sigma^{\prime}}$ (46)
$V_{j,\sigma;j^{\prime},\sigma^{\prime}}$ $=\langle c_{j^{\prime},\sigma^{\prime}}^{\dagger}(M-\Delta_{0})c_{j,\sigma}\rangle=\sum_{k}\bar{\mu}_{k}\langle c_{j^{\prime},\sigma^{\prime}}^{\dagger}\left[\left(n_{k,\uparrow}+n_{k,\downarrow}\right)-1\right]c_{j,\sigma}\rangle$
$=\sum_{k}\bar{\mu}_{k}\langle c_{j^{\prime},\sigma^{\prime}}^{\dagger}\left[n_{k,\uparrow}+n_{k,\downarrow},c_{j,\sigma}\right]\rangle=-\bar{\mu}_{j}\delta_{j,j^{\prime}}\left(\frac{1}{2}\tau_{0}+\overrightarrow{\tau}\cdot\langle\overrightarrow{S}_{j}\rangle\right)_{\sigma,\sigma^{\prime}}$ (47)
@@
$G^{1}_{j,\sigma;j^{\prime},\sigma^{\prime}}$ $=\delta_{j,j^{\prime}}\bar{\mu}_{j^{\prime}}\left(\tau_{0}\right)_{\sigma,\sigma^{\prime}}=M_{j,\sigma;j^{\prime},\sigma^{\prime}}$ (48)
We find that for $SU(2)$-symmetric on-site potentials, $G^{1}$ equals $M$. As second bare hopping Hamiltonian we investigate a typical spin-isotropic hopping WW($\equiv H_{0,l=1}$):
@@
The final expression for $G^{1}$ corresponding to the total bare Hamiltonian, $H_{0}=W+M$, is obtained by summing all individual contributions to $G^{1}$. Using $G^{0}$ and $G^{1}$, we can determine the Hamiltonian of zeros, $H^{\text{z}}$. For a spin-isotropic $H_{0}$ we find:
@@
As a final result, and in line with [33], we observe that the local terms of the bare Hamiltonian $M$ reemerge unrenormalized in the $H^{z}$-Hamiltonian, while the non-local terms $W$ are rescaled by the spin correlation function at the corresponding bond, $\langle\overrightarrow{S}_{j}\cdot\overrightarrow{S}_{j^{\prime}}\rangle_{j\neq j^{\prime}}$. The expectation values are taken with respect to the groundstate $|GS_{0}\rangle$ in the unperturbed limit, i.e., the groundstate of the corresponding effective spin-$1/2$ model (degenerated perturbation theory). Although $|GS_{0}\rangle$ is defined within the subspace of the atomic limit, it remains generally non-trivial, i.e., long-range correlations can occur:
@@
Due to the additive structure of the GF, Eq. (14), our results are valid for any density matrix $\rho=\sum_{\nu}w_{\nu}|E^{P,m}_{\nu,l}\rangle\langle E^{P,m}_{\nu,l}|$, which contains only SU(2)-symmetric states from the lowest energy sector m=0,l=0m=0,~l=0. In that case, all the expectation values can be simply updated:
@@
A.5 Green’s function approximation without SU(2)-symmetry
In the the previous section we explicitly assumed SU(2)-symmetry, what allows to truncate the perturbative series of the eigenstates and gain a rather simple expression on the GFZ-Hamiltonian. However, in order to derive the Green’s function approximation SU(2)-symmetry is not necessary. The Källen-Lehmann presentation of the Green’s function can be given as:
$G_{j,\sigma;j^{\prime},\sigma^{\prime}}(\omega)$ $=\sum_{\nu}\Bigg{[}\frac{\langle GS^{P}|c_{j,\sigma}|E_{\nu}^{P}\rangle\langle E_{\nu}^{P}|c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{P}\rangle}{\omega+E_{0}^{P}-E_{\nu}^{P}}+\frac{\langle E_{\nu}^{P}|c_{j,\sigma}|GS^{P}\rangle\langle GS^{P}|c^{\dagger}_{j^{\prime},\sigma^{\prime}}|E_{\nu}^{P}\rangle}{\omega+E_{\nu}^{P}-E_{0}^{P}}\Bigg{]}$ (58)
$=\mathcal{A}_{j,\sigma;j^{\prime},\sigma^{\prime}}+\mathcal{B}_{j,\sigma;j^{\prime},\sigma^{\prime}}$ (59)
for brevity we supress all the sub-indices, and sum over all eigenstates $\{|E_{\nu}\rangle\}$. In the following we will treat the GF perturbatively in $\phi$ and in the frequency $\omega$, both up to its first order. We send $\phi\rightarrow 0$, mathematically correct we thereby assume a fixed $U$ and $D\rightarrow 0$, this is however equivalent to the strongly correlated limit. Due to the similarity, we only give details on the expansion for the first term in Eq. (59). As first step, we apply the expansion in $\phi$:
$\mathcal{A}(\phi)\approx\mathcal{A}\big{|}_{\phi=0}+\phi\partial_{\phi}\mathcal{A}\big{|}_{\phi=0}$ (60)
The individual terms are:
$\mathcal{A}\big{|}_{\phi=0}$ $=\sum_{\nu}\frac{\langle GS^{0}|c_{j,\sigma}|E_{\nu}^{0}\rangle\langle E_{\nu}^{0}|c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{0}\rangle}{\omega+E_{0}^{0}-E_{\nu}^{0}}$ (61)
$\partial_{\phi}\mathcal{A}\big{|}_{\phi=0}$ $=\sum_{\nu}\bigg{[}\frac{\langle GS^{0}|c_{j,\sigma}\partial_{\phi}\left[|E_{\nu}^{P}\rangle\langle E_{\nu}^{P}|\right]\big{|}_{\phi=0}c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{0}\rangle}{\omega+E_{0}^{0}-E_{\nu}^{0}}+\frac{\langle E_{\nu}^{0}|c^{\dagger}_{j^{\prime},\sigma^{\prime}}\partial_{\phi}\left[|GS^{P}\rangle\langle GS^{P}|\right]\big{|}_{\phi=0}c_{j,\sigma}|E_{\nu}^{0}\rangle}{\omega+E_{0}^{0}-E_{\nu}^{0}}$ (62)
$-\frac{\langle GS^{0}|c_{j,\sigma}|E_{\nu}^{0}\rangle\langle E_{\nu}^{0}|c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{0}\rangle}{(\omega+E_{0}^{0}-E_{\nu}^{0})^{2}}\partial_{\phi}\left[E_{0}^{P}-E_{\nu}^{P}\right]\big{|}_{\phi=0}\bigg{]}$ (63)
In each term we recognize amplitudes like $\langle E_{\nu}^{0}|c^{(\dagger)}_{j^{\prime},\sigma^{\prime}}|GS^{0}\rangle$ which in combination with the unperturbed groundstate $|GS^{0}\rangle$ is non-vanishing only in the lowest energy sector (see Section A.1). Most important hereby is that the denominator becomes constant:
$\mathcal{A}\big{|}_{\phi=0}$ $=\sum_{\nu}\frac{\langle GS^{0}|c_{j,\sigma}|E_{\nu}^{0}\rangle\langle E_{\nu}^{0}|c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{0}\rangle}{\omega-U}$ (64)
$\partial_{\phi}\mathcal{A}\big{|}_{\phi=0}$ $=\sum_{\nu}\frac{\langle GS^{0}|c_{j,\sigma}\partial_{\phi}\left[|E_{\nu}^{P}\rangle\langle E_{\nu}^{P}|\right]\big{|}_{\phi=0}c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{0}\rangle}{\omega-U}+\frac{\langle E_{\nu}^{0}|c^{\dagger}_{j^{\prime},\sigma^{\prime}}\partial_{\phi}\left[|GS^{P}\rangle\langle GS^{P}|\right]\big{|}_{\phi=0}c_{j,\sigma}|E_{\nu}^{0}\rangle}{\omega-U}$ (65)
$-\frac{\langle GS^{0}|c_{j,\sigma}|E_{\nu}^{0}\rangle\langle E_{\nu}^{0}|c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{0}\rangle}{(\omega-U)^{2}}\partial_{\phi}\left[E_{0}^{P}-E_{\nu}^{P}\right]\Big{|}_{\phi=0}$ (66)
In the next step we want to take the sum over all eigenstates $\{|E_{\nu}\rangle\}$. Here, Eq. (64) can be easily evaluated:
$\mathcal{A}\big{|}_{\phi=0}$ $=\frac{\langle GS^{0}|c_{j,\sigma}c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{0}\rangle}{\omega-U}$ (67)
In Eq. (66), the first term will vanish, what can be seen by using that the summation and the derivative commutes:
$\sum_{\nu}\partial_{\phi}\left[|E_{\nu}^{P}\rangle\langle E_{\nu}^{P}|\right]=\partial_{\phi}\bigg{[}\sum_{\nu}|E_{\nu}^{P}\rangle\langle E_{\nu}^{P}|\bigg{]}=\partial_{\phi}\mathbb{I}=0\text{.}$ (68)
The second term can be easily evaluated:
$\sum_{\nu}\frac{\langle E_{\nu}^{0}|c^{\dagger}_{j^{\prime},\sigma^{\prime}}\partial_{\phi}\left[|GS^{P}\rangle\langle GS^{P}|\right]\big{|}_{\phi=0}c_{j,\sigma}|E_{\nu}^{0}\rangle}{\omega-U}=\partial_{\phi}\frac{\langle GS^{P}|c_{j,\sigma}c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{P}\rangle}{\omega-U}\big{|}_{\phi=0}\text{.}$ (69)
While the evaluation of the third term is more subtle and involves general results from degenerated perturbation theory [48]:
$D\partial_{\phi}\langle E_{\nu}^{P}|H_{H}|E_{\nu}^{P}\rangle\Big{|}_{\phi=0}$ $=U\langle E_{\nu}^{0}|H_{0}|E_{\nu}^{0}\rangle$ (70)
$H_{0}|E_{\nu,l}^{0,m}\rangle$ $=\langle E_{\nu,l}^{0,m}|H_{0}|E_{\nu,l}^{0,m}\rangle|E_{\nu,l}^{0,m}\rangle+\sum_{\nu^{\prime},l^{\prime},m^{\prime};l^{\prime}\neq l\wedge m^{\prime}\neq m}c_{\nu^{\prime},l^{\prime},m^{\prime}}|E_{\nu^{\prime},l^{\prime}}^{0,m^{\prime}}\rangle\text{.}$ (71)
Consequently we find:
- $\sum_{\nu}\frac{\langle GS^{0}|c_{j,\sigma}|E_{\nu}^{0}\rangle\langle E_{\nu}^{0}|c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{0}\rangle}{(\omega-U)^{2}}\partial_{\phi}\left[E_{0}^{P}-E_{\nu}^{P}\right]\Big{|}_{\phi=0}$ (72)
$=-U\frac{\langle GS^{0}|c_{j,\sigma}c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{0}\rangle\langle GS^{0}|H_{0}|GS^{0}\rangle}{D(\omega-U)^{2}}+U\frac{\langle GS^{0}|c_{j,\sigma}H_{0}c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{0}\rangle}{D(\omega-U)^{2}}\text{.}$ (73)
As the next step we take the expansion at low energy $\omega=0$ to the first order:
$\mathcal{A}(\phi,\omega)\approx\mathcal{A}\big{|}_{\phi=0,\omega=0}+\phi\partial_{\phi}\mathcal{A}\big{|}_{\phi=0,\omega=0}+\omega\partial_{\omega}\mathcal{A}\big{|}_{\phi=0,\omega=0}\text{.}$ (74)
The individual terms are:
$\mathcal{A}\big{|}_{\phi=0,\omega=0}$ $=-\frac{1}{U}\langle GS^{0}|c_{j,\sigma}c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{0}\rangle$ (75)
$\partial_{\phi}\mathcal{A}\big{|}_{\phi=0,\omega=0}$ $=-\frac{1}{UD}\langle GS^{0}|c_{j,\sigma}c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{0}\rangle\langle GS^{0}|H_{0}|GS^{0}\rangle+\frac{1}{UD}\langle GS^{0}|c_{j,\sigma}H_{0}c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{0}\rangle$ (76)
$\quad\quad-\frac{1}{U}\partial_{\phi}\langle GS^{P}|c_{j,\sigma}c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{P}\rangle\big{|}_{\phi=0}$ (77)
$\partial_{\omega}\mathcal{A}\big{|}_{\phi=0,\omega=0}$ $=-\frac{1}{U^{2}}\langle GS^{0}|c_{j,\sigma}c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{0}\rangle$ (78)
An equivalent derivation gives the result on $\mathcal{B}$, combining both results in the low energy Green’s function approximation leads to:
$G_{j,\sigma;j^{\prime},\sigma^{\prime}}(\omega)$ $\approx\frac{1}{U^{2}}\left[H^{\text{z}}_{j,\sigma;j^{\prime},\sigma^{\prime}}-\omega\delta_{j,\sigma;j^{\prime},\sigma^{\prime}}\right]$ (79)
$H^{\text{z}}_{j,\sigma;j^{\prime},\sigma^{\prime}}$ $=-U\langle GS^{0}|\left[c_{j,\sigma},c^{\dagger}_{j^{\prime},\sigma^{\prime}}\right]|GS^{0}\rangle-D\partial_{\phi}\langle GS^{P}|\left[c_{j,\sigma},c^{\dagger}_{j^{\prime},\sigma^{\prime}}\right]|GS^{P}\rangle\big{|}_{\phi=0}$ (80)
$-\langle GS^{0}|\left[c_{j,\sigma},c^{\dagger}_{j^{\prime},\sigma^{\prime}}\right]|GS^{0}\rangle\langle GS^{0}|H_{0}|GS^{0}\rangle+\langle GS^{0}|c_{j,\sigma}H_{0}c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{0}\rangle$ (81)
$-\langle GS^{0}|c^{\dagger}_{j^{\prime},\sigma^{\prime}}H_{0}c_{j,\sigma}|GS^{0}\rangle$ (82)
At this stage we already derived one of our main results of the manuscript. We only assumed that the Hubbard Hamiltonian at study is strongly correlated $\phi\ll 1$ and at half-filling. In comparison with Eq. (34) we recognize almost all terms, up to the derivative $\partial_{\phi}\langle GS^{P}|\left[c_{j,\sigma},c^{\dagger}_{j^{\prime},\sigma^{\prime}}\right]|GS^{P}\rangle\big{|}_{\phi=0}$, which can be treated by perturbation theory [51]. In the case of an SU(2)-symmetric total Hamiltonian and an SU(2)-symmetric ground state, our expansion in Eq. (13) is sufficient and yields the exact result. This, however, is by no means trivial; in fact, it follows from second-order degenerate perturbation theory in combination with the spin-operator representation Eq. (35) and the Wigner–Eckart theorem[48], which allows one to significantly simplify the expectation value $\left.\partial_{\phi}\langle GS^{P}|\left[c_{j,\sigma},c^{\dagger}_{j^{\prime},\sigma^{\prime}}\right]|GS^{P}\rangle\right|_{\phi=0}$.
Morover, the GFZ-Hamiltonian in Eq. (82) reveals already the structure of the GFZ-dispersion. In particular we recognize two different energy scales, $\mathcal{O}(U)$ and $\mathcal{O}(D)$. The dominant zeroth order contribution $2U\Big{[}\langle GS^{0}|\overrightarrow{S}_{j}|GS^{0}\rangle\cdot\overrightarrow{\tau}\Big{]}_{\sigma,\sigma^{\prime}}\delta_{j,j^{\prime}}$, is local and directly related to the magnetic polarization of the strongly correlated groundstate. It shifts the zeros towards the Hubbard bands, what represents the main effect of a magnetic field. The investigated fine-structure of the dispersion is created by the remaining terms of order $\mathcal{O}(D)$, its evaluation involves an exact treatment of the correlators, which can become very lengthy for the SU(2)-breaking states or/and Hamiltonians. Nonetheless, for the SU(2)-symmetric case we arrive at an very elegant results as we derive in Section A.3 and A.4.
A.6 Self-consistent selfenergy construction